# Classification of left invariant Riemannian metrics on nonunimodular 4-dimensional Lie groups

### v2

@@ 4.1 The Lie algebra $\operatorname{A}_{4,2}^{\alpha}$
The Lie algebra $\operatorname{A}_{4,2}^{\alpha}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$[e_{1},e_{4}]=\alpha e_{1},\quad[e_{2},e_{4}]=e_{2},\quad[e_{3},e_{4}]=e_{2}+e_{3}.$
@@ 4.1.1 The Lie algebra $\operatorname{A}_{4,2}^{\alpha}$ where $\alpha\neq(0,1)$
This matrix is orthogonal precesily when $b_{44}=c_{44}$, $a_{7}=0$, $a_{1},a_{6}$ are equal to ±$\pm 1$, $b_{22}=c_{22}$, $a_{1}b_{12}=a_{6}c_{12}$ and $a_{1}b_{13}=a_{6}c_{13}$. By choice of $a_{1}$ and $a_{6}$ we can assume that $b_{12}\geq 0$, but we have $b_{13}\in\mathbb{R}$. ∎
@@ 4.1.2 The Lie algebra $\operatorname{A}_{4,2}^{1}$
This matrix is orthogonal precesily when $a_{1},a_{6}$ are equal to ±$\pm 1$, $b_{44}=c_{44}$, $b_{22}=c_{22}$ and $a_{1}b_{12}=a_{6}c_{12}$. By choice of $a_{1}$ and $a_{6}$ we can put $b_{12}\geq 0$. ∎
@@ 4.2 The Lie algebra $\operatorname{A}_{4,3}$
The Lie algebra $\operatorname{A}_{4,3}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$[e_{1},e_{4}]=e_{1},\quad[e_{3},e_{4}]=e_{2}.$
@@ 4.3 The Lie algebra $\operatorname{A}_{4,4}$
The Lie algebra $\operatorname{A}_{4,4}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$[e_{1},e_{4}]=e_{1},\quad[e_{2},e_{4}]=e_{1}+e_{2},\quad[e_{3},e_{4}]=e_{2}+e_{3}.$
@@ 4.4 The Lie algebra $\operatorname{A}_{4,5}^{\alpha,\beta}$
The Lie algebra $\operatorname{A}_{4,5}^{\alpha,\beta}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$[e_{1},e_{4}]=e_{1},\quad[e_{2},e_{4}]=\alpha e_{2},\quad[e_{3},e_{4}]=\beta e_{3}.$
@@
The matrix $C^{-1}AB$ is orthogonal precisely when $b_{44}=c_{44}$, $a_{1},a_{6}$ and $a_{11}$ are equal to ±$\pm 1$ and $x=y=z=0$. Thus we obtain that
@@ 4.5 The Lie algebra $\operatorname{A}_{4,6}^{\alpha,\beta}$
The Lie algebra $\operatorname{A}_{4,6}^{\alpha,\beta}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$[e_{1},e_{4}]=\alpha e_{1},\quad[e_{2},e_{4}]=\beta e_{2}-e_{3},\quad[e_{3},e_{4}]=e_{2}+\beta e_{3}.$
@@
If $10<b_{33}<$ 1, then the condition that the middle $2\times 2$ block of $C^{-1}AB$ is orthogonal implies that $a_{7}=0$. Thus $\begin{bmatrix}b_{12}\\ b_{13}\end{bmatrix}=\pm\begin{bmatrix}c_{12}\\ c_{13}\end{bmatrix}.$
@@ 4.6 The Lie algebra $\operatorname{A}_{4,7}$
The Lie algebra $\operatorname{A}_{4,7}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$[e_{2},e_{3}]=e_{1},\quad[e_{1},e_{4}]=2e_{2},\quad[e_{2},e_{4}]=e_{2},\quad[e_{3},e_{4}]=e_{2}+e_{3}.$
@@ 4.7 The Lie algebra $\operatorname{A}_{4,9}^{\beta}$
The Lie algebra $\operatorname{A}_{4,9}^{\beta}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$[e_{2},e_{3}]=e_{1},\quad[e_{1},e_{4}]=(1+\beta)e_{1},\quad[e_{2},e_{4}]=e_{2},\quad[e_{3},e_{4}]=\beta e_{3}.$
@@
This matrix is orthogonal precesily when $a_{6}$, $a_{11}$ are equal to ±$\pm 1$, $b_{11}=c_{11}$, $b_{44}=c_{44}$ and $x=y=z=0$. This means that
@@ 4.8 The Lie algebra $\operatorname{A}_{4,11}^{\alpha}$
The Lie algebra $\operatorname{A}_{4,11}^{\alpha}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$[e_{2},e_{3}]=e_{1},\quad[e_{1},e_{4}]=2\alpha e_{1},\quad[e_{2},e_{4}]=\alpha e_{2}-e_{3},\quad[e_{3},e_{4}]=e_{2}+\alpha e_{3}.$
@@ 4.9 The Lie algebra $\operatorname{A}_{4,12}$
The Lie algebra $\operatorname{A}_{4,12}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the only nonzero brackets are
$[e_{1},e_{3}]=e_{1},\quad[e_{2},e_{3}]=e_{2},\quad[e_{1},e_{4}]=-e_{2},\quad[e_{2},e_{4}]=e_{1}.$
@@
If $10<b_{22}<$ 1, then the condition that the upper $2\times 2$ block of $C^{-1}AB$ is orthogonal implies that $a_{2}=0$. In this case we have

### v3

@@ 4.1 The Lie algebra $\operatorname{A}_{4,2}^{\alpha}$
The Lie algebra $\operatorname{A}_{4,2}^{\alpha}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$e_{1},e_{4}]=\alpha e_{1},\quad[e_{2},e_{4}]=e_{2},\quad[e_{3},e_{4}]=e_{2}+e_{3}.$
@@ 4.1.1 The Lie algebra $\operatorname{A}_{4,2}^{\alpha}$ where $\alpha\neq(0,1)$
This matrix is orthogonal precesily when $b_{44}=c_{44}$, $a_{7}=0$, $a_{1},a_{6}$ are equal to $\pm 1$, $b_{22}=c_{22}$, $a_{1}b_{12}=a_{6}c_{12}$ and $a_{1}b_{13}=a_{6}c_{13}$. By choice of $a_{1}$ and $a_{6}$ we can assume that $b_{12}\geq 0$, but we have $b_{13}\in\mathbb{R}$. ∎
@@ 4.1.2 The Lie algebra $\operatorname{A}_{4,2}^{1}$
This matrix is orthogonal precesily when $a_{1},a_{6}$ are equal to $\pm 1$, $b_{44}=c_{44}$, $b_{22}=c_{22}$ and $a_{1}b_{12}=a_{6}c_{12}$. By choice of $a_{1}$ and $a_{6}$ we can put $b_{12}\geq 0$. ∎
@@ 4.2 The Lie algebra $\operatorname{A}_{4,3}$
The Lie algebra $\operatorname{A}_{4,3}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$e_{1},e_{4}]=e_{1},\quad[e_{3},e_{4}]=e_{2}.$
@@ 4.3 The Lie algebra $\operatorname{A}_{4,4}$
The Lie algebra $\operatorname{A}_{4,4}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$e_{1},e_{4}]=e_{1},\quad[e_{2},e_{4}]=e_{1}+e_{2},\quad[e_{3},e_{4}]=e_{2}+e_{3}.$
@@ 4.4 The Lie algebra $\operatorname{A}_{4,5}^{\alpha,\beta}$
The Lie algebra $\operatorname{A}_{4,5}^{\alpha,\beta}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$e_{1},e_{4}]=e_{1},\quad[e_{2},e_{4}]=\alpha e_{2},\quad[e_{3},e_{4}]=\beta e_{3}.$
@@
The matrix $C^{-1}AB$ is orthogonal precisely when $b_{44}=c_{44}$, $a_{1},a_{6}$ and $a_{11}$ are equal to $\pm 1$ and $x=y=z=0$. Thus we obtain that
@@ 4.5 The Lie algebra $\operatorname{A}_{4,6}^{\alpha,\beta}$
The Lie algebra $\operatorname{A}_{4,6}^{\alpha,\beta}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$e_{1},e_{4}]=\alpha e_{1},\quad[e_{2},e_{4}]=\beta e_{2}-e_{3},\quad[e_{3},e_{4}]=e_{2}+\beta e_{3}.$
@@
If $0<b_{33}<1$, then the condition that the middle $2\times 2$ block of $C^{-1}AB$ is orthogonal implies that $a_{7}=0$. Thus $\begin{bmatrix}b_{12}\\ b_{13}\end{bmatrix}=\pm\begin{bmatrix}c_{12}\\ c_{13}\end{bmatrix}$.
@@ 4.6 The Lie algebra $\operatorname{A}_{4,7}$
The Lie algebra $\operatorname{A}_{4,7}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$e_{2},e_{3}]=e_{1},\quad[e_{1},e_{4}]=2e_{2},\quad[e_{2},e_{4}]=e_{2},\quad[e_{3},e_{4}]=e_{2}+e_{3}.$
@@ 4.7 The Lie algebra $\operatorname{A}_{4,9}^{\beta}$
The Lie algebra $\operatorname{A}_{4,9}^{\beta}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$e_{2},e_{3}]=e_{1},\quad[e_{1},e_{4}]=(1+\beta)e_{1},\quad[e_{2},e_{4}]=e_{2},\quad[e_{3},e_{4}]=\beta e_{3}.$
@@
This matrix is orthogonal precesily when $a_{6}$, $a_{11}$ are equal to $\pm 1$, $b_{11}=c_{11}$, $b_{44}=c_{44}$ and $x=y=z=0$. This means that
@@ 4.8 The Lie algebra $\operatorname{A}_{4,11}^{\alpha}$
The Lie algebra $\operatorname{A}_{4,11}^{\alpha}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$e_{2},e_{3}]=e_{1},\quad[e_{1},e_{4}]=2\alpha e_{1},\quad[e_{2},e_{4}]=\alpha e_{2}-e_{3},\quad[e_{3},e_{4}]=e_{2}+\alpha e_{3}.$
@@ 4.9 The Lie algebra $\operatorname{A}_{4,12}$
The Lie algebra $\operatorname{A}_{4,12}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the only nonzero brackets are
$e_{1},e_{3}]=e_{1},\quad[e_{2},e_{3}]=e_{2},\quad[e_{1},e_{4}]=-e_{2},\quad[e_{2},e_{4}]=e_{1}.$
@@
If $0<b_{22}<1$, then the condition that the upper $2\times 2$ block of $C^{-1}AB$ is orthogonal implies that $a_{2}=0$. In this case we have